OpenStudy (anonymous):

find an equation of the line having the given slope and containing the given point.

OpenStudy (anonymous):

m = 4/5, (7,6)

OpenStudy (anonymous):

for this you want to use the point-slope form of an equation: \[y-y_1 =m(x-x_1)\] where m is your slope, and x_1 and y_1 are your point.

OpenStudy (anonymous):

y = mx + b ==> 6 = (4/5)(7) + b find the value of b. the equation is y = (4/5)x + [value of b] m and b must always be defined, whereas x and y are variables.